# AG-SLAM: Active Gaussian Splatting SLAM

### v1

@@ 3.2 Information-driven Path Proposal
We consider an agent moving in a 2D plane (e.g. a wheeled robot), so we construct a 2D occupancy map for path planning by ground projecting the means of 3D Gaussians.
Frontier-based exploration [58] is incorporated into our algorithm to provide candidates for view selection. The frontiers are defined to be points on the boundary between free space and unobserved space (not marked as free or occupied) in our occupancy map. We then form an initial set of candidate poses $\mathcal{T}_{I}$ by sampling poses around each point in the largest frontier. If there are no frontiers, we instead sample poses around the center of each Gaussian to form $\mathcal{T}_{I}$. We then evaluate the Expected Information Gain (EIG) for each pose ${\bf x}^{acq}_{i}\in\mathcal{T}_{I}$, given by
@@
$\mathcal{I}({\bf w})=\mathbb{E}_{{\bf x}\sim p({\bf x})}\left[{\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]\right]\simeq\sum_{k=1}^{N}{\mathbf{H}}^{\prime\prime}[{\bf y}_{k}|{\bf x}_{k},{\bf w}],\quad{\bf x}_{k}\sim p({\bf x})$ (10)
where $p({\bf x}_{k})$ is approximated with a uniform distribution of camera poses in the free space of the current map. Besides, we also uniformly initialize 3D Gaussians in the space, which will be subsequently updated with rendering losses for visited regions. In Fig. 2, we show the relationship between PSNR and EIG on sampled poses in the Cantwell scene. The result aligns with common sense that EIG should decrease as PSNR increases. We also show some cases on the lower-left and upper-right of the scatter plot. For the lower left capture, even though the left part is poorly reconstructed, most of view is occupied by textureless wall, leading to low EIG score. For the upper right capture, even though the scene has a moderate reconstruction, the content is rich so our algorithm returns a high EIG score.
Finally, we compute paths towards each pose in $\mathcal{T}_{F}$ with the A* algorithm [13] using the occupancy map, selecting which path to follow as described in Section 3.3. The path can be defined as an ordered set of camera poses from the current location $x_{t}$ at exploration step $t$ to the frontier points $x_{T}^{j}$.
@@
3.3 Path Selection with Localization Uncertainty
Following FisherRF [19], the EIG for 3D Gaussian parameters along a path $\mathcal{P}_{j}$ can be computed as:
$\sum_{{\bf x}_{i}\in\mathcal{P}_{j}}\mathtt{EIG}_{\mathcal{P}_{j},i}({\bf x}_{i}),\;\;\mathtt{EIG}_{\mathcal{P}_{j},i}({\bf x}_{i})=\operatorname{tr}\left({\mathbf{H}}^{\prime\prime}[{\bf y}_{i}|{\bf x}_{i},{\bf w}]\;\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})^{-1}\right)$ (12)
[…]
$\mathrm{Cov}(T(\hat{\tau_{i}}))\geq\mathcal{I}(\tau_{i})^{-1}$ (14)
Matsuki et al. [30] computed the Jacobians of camera pose with respect to the mean and covariances of each gaussian $\frac{\partial{\bm{\mu}_{I}}}{\partial{{\bf x}}}$ and $\frac{\partial{\bm{\Sigma}_{I}}}{\partial{{\bf x}}}$. However, we need to compute the Jacobian of $\tau_{i}$ with respect to the rendering output:
$\nabla_{\tau_{i}}f(\tau_{i},{\bf w})=\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\tau_{i}}}=\begin{bmatrix}\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\bm{\mu}_{I}}}&\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\bm{\Sigma}_{I}}}\end{bmatrix}\begin{bmatrix}\frac{\mathcal{D}{\bm{\mu}_{C}}}{\mathcal{D}{\tau_{i}}}\\ \frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{\tau_{i}}}\end{bmatrix}$ (16)
where $\eta$ is a hyper-parameter controlling the importance between EIG and localization accuracy. The agent can then explore the environment with planned path $\mathcal{P}$. Our active SLAM system constantly updates the map, and we replan using our active path planning algorithm if we detect the agent is getting close to a possible obstacle or upon reaching the end of the previously selected path.
4 Experiments
4.1 Experimental set-up
Dataset
Our algorithm is evaluated in the Habitat Simulator [51] on the Gibson [57] and Habitat-Matterport 3D (HM3D) [39] datasets, which are comprised of indoor scenes reconstructed from scans of real houses. For Gibson we use all the scenes in the val split. For HM3D we use 5 scenes from the train split. We adopt the default start point in the Habitat Simulator as the starting point for exploration in each scene. The total number of steps for each experiment is 2000. The system takes color and depth images at the resolution of 800x800 and outputs a discrete action at each step. The action space consists of MOVE FORWARD by 5cm, TURN LEFT, and TURN RIGHT by 5°$\degree$. The field of view (FOV) is set to 90°$\degree$ vertically and horizontally. Please refer to the appendix for more details about the evaluation split and other hyper-parameters.
Metrics
[…]
Baselines
We compare to two exploration methods which assume ground truth pose: UPEN [11] and Active Neural Mapping (active-INR) [61]. UPEN uses an ensemble of models to predict and estimate the epistemic uncertainty of the occupancy map outside the field of view and leverages these to construct paths that reduce the uncertainty of the occupancy prediction. We also report the results of UPEN(gt), which uses ground truth pose as a reference because we find UPEN failed on some scenes due to localization failure. Active-INR aims to minimize the neural variability, that is the prediction robustness against random weight perturbation, of its signed distance field scene representation. We also compare our method with Active Neural SLAM (ANS) [4], explORB [36] and Frontier Based Exploration (FBE) [58] without ground truth pose provided. ANS learns to predict a map and pose estimate and a global goal based on them, which is then reached using a combination of a classical path planner and a trained local policy. ExplORB [36] adds possible loop closure based on the co-visibility of future pose and existing poses based on existing landmarks. It computes the Fisher Information of the Hessian on the pose graph optimization. FBE can be considered an ablation of our method to validate the importance of considering the scene and localization uncertainty, where instead of choosing paths using Eq. 17 we instead select the frontier based on the ratio of its area to the agent’s distance from it.
To make a fair comparison of the rendering quality, we run all the baselines using the MonoGS [30] backend for reconstruction. We run UPEN and FBE online but for ANS, active-INR and ExplORB we record and play back trajectories obtained using their codebase. Because the forward step size for ANS is much larger than for our method, we interpolate the trajectory so that the forward step size matches that of our method to make the steps comparable. For ExplORB, since the official implementation is based on MoveBase, which uses velocity commands, we sample the trajectory at 5 Hz. This will limit the distance change between frames to around 4cm and ${5}^{\circ}$ for the linear and angular distance, respectively, with the maximal linear and angular velocities as 20 cm/s and 0.5 rad/s, respectively. We also found that ANS and active-INR failed on some scenes due to localization failure. ANS produces a pose estimate (using information from noisy pose sensors not provided to our pipeline), so we set the pose estimate of the MonoGS backend to the one from ANS. As active-INR does not produce a pose estimate we evaluate it using the ground-truth pose.
4.2 Comparison Against Previous Methods
Table 1 shows the results of our method AG-SLAM and the baselines for exploration in scenes from the Gibson dataset, and Table 2 shows the results of our method and some baselines on HM3D. Note that the percentage coverage reported is the average percentage coverage per scene. As the scenes are different sizes this means that a method can have a lower percentage coverage with a higher area coverage, for example if it has better coverage in larger scenes but worse in smaller scenes. Our AG-SLAM outperforms the baselines on all metrics.
ANS [4] Active-INR [61] UPEN [11]
Refer to caption Refer to caption Refer to caption
ExplORB [36] FBE [58] AG-SLAM (Ours)
Refer to caption Refer to caption Refer to caption
ANS [4] Active-INR [61] UPEN [11]
Refer to caption Refer to caption Refer to caption
ExplORB [36] FBE [58] AG-SLAM (Ours)
Refer to caption Refer to caption Refer to caption
@@ 4.2 Comparison Against Previous Methods
Table 1: Evaluation of AG-SLAM and baselines on scenes from the Gibson dataset
@@
ANS 16.34 0.6818 0.3923 0.3886 0.1105 10.49 90.50
Active-INR (gt) 22.66 0.7652 0.2164 0.1528 - 9.20 78.34
UPEN (gt) 21.31 0.7325 0.2714 0.1696 - 8.79 76.17
UPEN 16.44 0.6678 0.4134 0.4841 0.5158 8.58 75.82
ExplORB 18.99 0.7175 0.3994 0.2664 0.2296 9.00 76.83
FBE 21.45 0.7618 0.2126 0.1028 0.0168 10.64 85.81
AG-SLAM (ours) 23.08 0.7959 0.1794 0.0763 0.0155 11.23 91.30
Table 2: Evaluation of AG-SLAM and baselines on scenes from the HM3D dataset
[…]
UPEN (gt) 15.58 0.5175 0.3936 0.4548 - 12.69 53.78
UPEN 12.23 0.4795 0.5157 0.7356 0.4393 10.72 44.96
FBE 15.80 0.5952 0.4392 0.4085 1.2004 15.69 66.44
AG-SLAM (ours) 18.74 0.6277 0.3534 0.1757 0.0208 19.70 80.96
@@
We further qualitatively compare the reconstruction qualities after active exploration in Fig. 3 and Fig. 4, and the trajectories in Fig. 5. We can see that AG-SLAM does not have major errors from failed localization and we have fewer gaps in the scenes than other methods. For example, in the Ribera scene all methods except for us and FBE miss the bathroom in the bottom left, and FBE misses more of the area around the sofa than us. For the trajectories, we show the estimated and ground truth trajectories for the Cantwell scene from the Gibson dataset. Cantwell is a relatively large and challenging scene, so it is suitable for showing the differences between methods. We show only a few baselines to keep the figure legible. We can see that Active-INR stays in a smaller area than the rest of the methods. ANS often goes close to walls, whereas FBE and AG-SLAM are generally more towards the center of corridors or rooms. ANS also extends further into rooms than these methods. FBE does not go to the region on the bottom right and ANS does not go to the region in the bottom left, whereas our AG-SLAM efficiently visits the main areas of the scene.
@@ 4.3 Ablation Study
To validate the effectiveness of using localization uncertainties, we performed an ablative study of our SLAM system with and without localization uncertainty for the Gibson dataset in Table. 3. As can be seen, the average trajectory error is also much lower with the localization term than without. In addition, the model performs consistently better for rendering when considering the localization uncertainty, as localization accuracy plays a key role in the optimization of Gaussian Splatting representation, and the quality of rendering is also dependent on localization accuracy. The percentage coverage is slightly better when not including the localization term; this can be explained by the method focusing more on exploration when it no longer balances it with localization accuracy. The area coverage is higher with the localization term, however, indicating that better localization could benefit coverage in larger scale scenes.
Table 3: Ablation Study of Localization Uncertainty Term on Scenes from the Gibson Dataset. We compare our method with and without the localization uncertainty term to validate that including it provides improvements on both localization and reconstruction
[…]
w.o. Localization Uncertainty 22.35 0.7830 0.3089 0.0823 0.1890 10.85 91.75
AG-SLAM 23.08 0.7959 0.1794 0.0763 0.0155 11.23 91.30
@@ 5 Conclusion
Recent SLAM methods employ the 3D Gaussian Splatting (3DGS) representation of the world, enabling a volumetric rendering as a measurement prediction. In this paper, we introduced active pose selection for 3DGS-based SLAM. Our AG-SLAM balances the information gain with respect to both location and the map. We mathematically formulated the expected information gain using the Fisher Information matrix and the Cramer-Rao Lower Bound. We evaluate our method for active SLAM on scenes from the Gibson [57] and Habitat-Matterport 3D  [39] datasets, in terms of the rendering quality, coverage and average tracking error. We show our uncertainty-based criteria for path selection improves over using frontier-based exploration [58] with a selection criteria that uses the frontier area and distance. We also compare our method with four recent state-of-the-art methods and show that AG-SLAM has superior performance.
To enable AG-SLAM to support more robotics applications, future work could extend AG-SLAM to consider movement with higher degrees of freedom (DOF) than the currently supported 3DOF. Incorporating semantic features [67; 46; 38] to allow for grounding language to the scene would also enable many robotics and computer vision applications.

### v2

@@ 3.3 Information-driven Path Proposal
After identifying a long horizon with the multimodal LLM, we form an initial set of candidate poses $\mathcal{T}_{I}$ by sampling points in the coarse region of interest from the LLM. If there are no unvisited boundaries or the multimodal LLM suggests we do not need to explore the frontiers, we sample poses across the free space to form $\mathcal{T}_{I}$. We then evaluate the Expected Information Gain (EIG) for each pose ${\bf x}^{acq}_{i}\in\mathcal{T}_{I}$, given by
@@
$\mathcal{I}({\bf w})=\mathbb{E}_{{\bf x}\sim p({\bf x})}\left[{\mathbf{H}}^{\prime\prime}[{\bf y}|{\bf x},{\bf w}]\right]\simeq\sum_{k=1}^{N}{\mathbf{H}}^{\prime\prime}[{\bf y}_{k}|{\bf x}_{k},{\bf w}],$ (6)
where ${\bf x}_{k}$ is drawn from a uniform distribution of camera poses in the free space of the current map which we use to approximate $p$. Besides, we also uniformly initialize 3D Gaussians in the space, which will be subsequently updated with rendering losses for visited regions. In Fig. 2, we show the relationship between PSNR and EIG on sampled poses in the Cantwell scene. The result aligns with our expectation that EIG should decrease as PSNR increases. We also show some cases on the lower-left and upper-right of the scatter plot. For the lower left capture, even though the left part is poorly reconstructed, most of the view is occupied by a textureless wall, leading to a low EIG. For the upper right capture, even though the scene has a moderate reconstruction, the content is rich, so our algorithm returns a high EIG.
Finally, we compute paths towards each pose in $\mathcal{T}_{F}$ with the A* algorithm [24] using the occupancy map, selecting which path to follow as described in Section 3.4. The path can be defined as an ordered set of camera poses from the current location $x_{t}$ at exploration step $t$ to the frontier points $x_{T}^{j}$ from the long horizon planning.
@@
ANS [7] Active-INR [85] UPEN [22] ExplORB [53] FBE [82] Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
3.4 Short-term Path Selection with Localization Uncertainty
The EIG for 3D Gaussian parameters along a path $\mathcal{P}_{j}$ can be computed as the sum over the path of the following term [31]:
$\mathtt{EIG}_{\mathcal{P}_{j},i}({\bf x}_{i})=\operatorname{tr}\left({\mathbf{H}}^{\prime\prime}[{\bf y}_{i}|{\bf x}_{i},{\bf w}]\;\mathcal{I}_{\mathcal{P}_{j},i}({\bf w})^{-1}\right)$ (8)
@@ 3.4 Short-term Path Selection with Localization Uncertainty
where ${\mathbf{H}}^{\prime\prime}[{\bf x}]$ is short for $-\nabla^{2}\log p(w)$ for clarity. If solely maximizing the $\mathtt{EIG}$, the robot will be more likely to explore unvisited regions. However, exploring regions that have not been well reconstructed also means the agent would have the risk of worse localization accuracy due to noise and ambiguities in the unreconstructed regions during pose optimization. The cost of localization must thus be considered during path planning to balance the importance of exploring new environments with maintaining localization accuracy. We propose to use Fisher Information as a measurement for the localization uncertainty that is also necessary for effective path planning for active Mapping algorithms. During optimization, we essentially optimize on the logarithmic mapping of $\tau_{i}\triangleq\text{Log}({\bf x}_{i})$ of our camera pose. By the Cramér–Rao bound, the covariance of $\tau_{i}\in\mathfrak{se}(3)$ can be lower-bounded with the inverse of Fisher Information matrix $\mathcal{I}(\tau_{i})$:
$\mathrm{Cov}(T(\hat{\tau_{i}}))\geq\mathcal{I}(\tau_{i})^{-1}$ (10)
Matsuki et al. [44] computed the Jacobians of the camera pose with respect to the mean and covariances of each gaussian $\frac{\partial{\boldsymbol{\mu}_{I}}}{\partial{{\bf x}}}$ and $\frac{\partial{\boldsymbol{\Sigma}_{I}}}{\partial{{\bf x}}}$. However, we need to compute the Jacobian of $\tau_{i}$ with respect to the rendering output:
$\nabla_{\tau_{i}}f(\tau_{i},{\bf w})=\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\tau_{i}}}=\begin{bmatrix}\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\boldsymbol{\mu}_{I}}}&\frac{\partial{f(\tau_{i},{\bf w})}}{\partial{\boldsymbol{\Sigma}_{I}}}\end{bmatrix}\begin{bmatrix}\frac{\mathcal{D}{\boldsymbol{\mu}_{C}}}{\mathcal{D}{\tau_{i}}}\\ \frac{\mathcal{D}{\mathbf{W}}}{\mathcal{D}{\tau_{i}}}\end{bmatrix}$ (12)
Unlike 3DGS parameters, our computation of the localization cost can be directly computed on the full Fisher Information matrix without using the Laplace approximation [12, 42]. That is because the full Jacobian $\nabla_{\tau_{i}}f(\tau_{i},{\bf w})$ is linear to the number of Gaussians and can be efficiently computed with our CUDA implementation.
where $\eta$ is a hyper-parameter controlling the importance between EIG and localization accuracy. The agent can then explore the environment with planned path $\mathcal{P}$. Our active mapping system constantly updates the map, and we replan using our active path planning algorithm if we detect the agent is getting close to a possible obstacle or upon reaching the end of the previously selected path.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
ANS 16.34 0.6818 0.3923 0.3886 0.1105 10.49
Active-INR (gt) 22.66 0.7652 0.2164 0.1528 - 9.20
UPEN (gt) 21.31 0.7325 0.2714 0.1696 - 8.79
UPEN 16.44 0.6678 0.4134 0.4841 0.5158 8.58
ExplORB 18.99 0.7175 0.3994 0.2664 0.2296 9.00
FBE 21.45 0.7618 0.2126 0.1028 0.1680 10.64
Ours 23.28 0.8067 0.2507 0.0696 0.0226 10.88
Table 1: Quantitative Evaluation on Reconstruction Quality and Tracking Accuracy on Gibson dataset
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
UPEN (gt) 15.58 0.5175 0.3936 0.4548 - 12.69
UPEN 12.23 0.4795 0.5157 0.7356 0.4393 10.72
FBE 15.80 0.5952 0.4392 0.4085 1.2004 15.69
Ours 19.86 0.7127 0.4122 0.1666 0.0336 18.48
Table 2: Quantitative Evaluation on Reconstruction Quality and Tracking Accuracy on HM3D dataset
@@
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth MAE $\downarrow$ RMSE ATE $\downarrow$ Coverage ($m^{2}$) $\uparrow$
w.o. LLM & Localization Uncertainty 16.15 0.6550 0.6193 0.3409 0.2478 16.06
w.o. LLM 16.94 0.6799 0.5847 0.2887 0.1694 17.26
Ours 19.86 0.7127 0.4122 0.1666 0.0336 18.48
Table 3: Ablation Study of Localization Uncertainty Term on Scenes from the HM3D Dataset. We compare our method with and without the localization uncertainty term to validate that including it provides improvements on both localization and reconstruction
4 Experiments
4.1 Experimental Set-up
Dataset
Following previous methods [22, 85], our algorithm is evaluated in the Habitat Simulator [73] on the Gibson [81] and Habitat-Matterport 3D (HM3D) [56] datasets comprised of indoor scenes reconstructed from scans of real houses. We use all the scenes in the val split for the Gibson dataset. We also evaluate our method on single-floor scenes whose scene ID can be found in the supplementary. We adopt the default start point in the Habitat Simulator as the starting point for exploration in each scene. The total number of steps for each experiment is 2000. The system takes color and depth images at the resolution of 800x800 and outputs a discrete action at each step. The action space consists of MOVE FORWARD by 5cm, TURN LEFT, and TURN RIGHT by 5°$\degree$. The field of view (FOV) is set to 90°$\degree$ vertically and horizontally. Please refer to the supplement for more details about the evaluation split and other hyper-parameters.
Metrics
We evaluated our method using the Peak-signal-to-noise ratio (PSNR), Structural Similarity Index Measure  (SSIM) [79], Learned Perceptual Image Patch Similarity (LPIPS) [89] for RGB rendering and mean absolute error (MAE) for depth rendering as metrics for scene reconstruction quality. We calculate these metrics using 2000 points uniformly sampled from the movement plane of the agent in the scene, discarding any points that are not navigable. We argue that the rendering quality reflects both reconstruction quality and pose accuracy because high tracking accuracy would help the training of the 3D Gaussian Splatting model. Meanwhile, misaligned poses will lead to misaligned rendering at test time thus leading to inferior results. Following previous approaches [22, 7], we also use coverage in $m^{2}$ as an evaluation metric. To evaluate the pose estimation accuracy, we use the root mean squared average tracking error (RMSE ATE), but as the trajectories for each method are different, the RMSE ATE should only be considered along with other metrics such as coverage.
Baselines
We compare to two exploration methods which assume ground truth pose: UPEN [22] and Active Neural Mapping (active-INR) [85]. UPEN chooses trajectories to reduce the epistemic uncertainty of the occupancy prediction produced by an ensemble of models. Active-INR aims to minimize the prediction robustness against random weight perturbation of its signed distance field scene representation. We also compare our method with Active Neural SLAM (ANS) [7], explORB [53] and Frontier Based Exploration (FBE) [82] without ground truth pose provided. ANS [7] learns to predict a map and estimate a global goal based on it. ExplORB [53] computes the Fisher Information of the Hessian on its pose graph optimization.
To compare the rendering quality fairly, we run all the baselines using the MonoGS [44] backend for reconstruction. We run UPEN and FBE online, but for ANS, active-INR, and ExplORB, we record and playback trajectories obtained using their source code. Because the forward step size for ANS is much larger than for our method, we interpolate the trajectory so that the forward step size matches our method’s to make the steps comparable. For ExplORB, since the official implementation is based on MoveBase, which uses velocity commands, we sample the trajectory at 5 Hz. We also found that ANS, active-INR, and UPEN failed on some scenes due to localization failure. ANS produces a pose estimate (using information from noisy pose sensors not provided to our pipeline), so we set the pose estimate of the MonoGS backend to the one from ANS. As active-INR and UPEN do not produce a pose estimate, we evaluate them using the ground-truth pose.
4.2 Comparison Against Previous Methods
@@
Table 1 shows the results of our method and the baselines for exploration in scenes from the Gibson dataset [81], and Table 2 shows the results of our method and some baselines on HM3D [6]. Our method outperforms the baselines on all metrics. We further qualitatively compare the reconstruction qualities after active exploration in Fig. 4 and Fig. 5, and the trajectories in Fig. 6. Our method does not have major errors from failed localization, and we have fewer gaps in the scenes than other methods. For example, in the Ribera scene, all methods except for us and FBE miss the bathroom at the bottom left, and FBE misses more of the area around the sofa than us. For the trajectories, we show the estimated and ground truth trajectories for the Cantwell scene from the Gibson dataset. Cantwell is a relatively large and challenging scene, so it is suitable for showing the differences between methods. We show only a few baselines to keep the figure legible. We can see that Active-INR stays in a smaller area than the rest of the methods. ANS often goes close to walls, whereas FBE and our method are generally more toward the center of corridors or rooms. ANS also extends further into rooms than these methods. FBE does not go to the region on the bottom right, and ANS does not go to the region on the bottom left, whereas our method efficiently visits most areas of the scene.
@@ 4.3 Ablation Study
To validate the effectiveness of localization uncertainty and LLM, we performed an ablative study of each component in our system on the Gibson dataset in Table. 3. As can be seen, the average trajectory error is much lower with the localization uncertainty than without. In addition, after adding LLM, the model performs better in the overall rendering. The ATE slightly increases after adding LLM, but this could be attributed to the larger area traversed by the robot.
@@ 5 Conclusion
We present an active mapping system that can autonomously explore an environment, the first method to do so using a 3D Gaussian representation without ground truth camera poses. By rendering maps using our scene representation, we can seamlessly query a multimodal Large Language Model for long-horizon planning to leverage its knowledge of scene layouts. This allows the robot to efficiently explore the scene while avoiding setting unreachable navigation goals. We then plan an optimal path to the long-horizon goal using our information-driven path proposal and selection algorithm, which balances the information gain with respect to the map against the localization uncertainty, ensuring that the robot captures the geometric details of the scene for a high-quality reconstruction. We evaluate our method on scenes from the Gibson [81] and Habitat-Matterport 3D  [56] datasets in terms of the rendering quality, coverage, and average tracking error. We compare to recent exploration methods with [22, 85] and without [53, 7] ground-truth pose, using the same scene representation for reconstruction as us for fairness, and show that our method has superior performance. To enable our method to support more robotics applications, future work could extend it to take advantage of our 3D scene representation and consider movement with higher degrees of freedom (DOF) than the currently supported 3DOF. Incorporating semantic features [91, 68, 55] to allow for grounding language to the scene would also enable many robotics and computer vision applications.